# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 12 Main Case 1
$t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}c_{1}$,
$\sigma(t_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(c_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}c$ for some $c\in C\langle E,F\rangle$, where the rewrite rule $\sigma(c_{1})\dot{\rightarrow}c$ is in $R\langle E,F\rangle$, and
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma(c_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[c]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Case 2: $m=0$. Then we consider the rewrite rule $\sigma\dot{\rightarrow}b\in R\langle{E\cup F},\emptyset\rangle$. Then we have
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[b]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
@@
$\sigma\mathop{\rightarrow}_{R\langle E,F\rangle}c$ for some $c\in C\langle E,F\rangle$, where the rewrite rule $\sigma\dot{\rightarrow}c$ is in $R\langle E,F\rangle$, and
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta[c]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
()($\Leftarrow$)( ⇐ ) Let $t\in W$ and $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Then $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Consequently, $R\langle E,F\rangle$ or $R\langle F,E\rangle$ simulate $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Assume that $R\langle E,F\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. The case $R\langle F,E\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ is similar. By Lemma 12.104, there exists $d\in C\langle E,F\rangle$ with $d\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}d$. By Proposition 4.38, $d=t/_{\Theta\langle E,F\rangle}$. For this reason $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Consequently, for any $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. As a consequence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$, and hence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. ∎
@@ 12.1 Nested Partial Depth First Search on the Auxiliary dpwpa $AUX[E;F]$ for the GTESs $E$ and $F$
In this way, algorithm NPDFS decides the following formula: For each $a\in C\langle E\cup F,\emptyset\rangle$,
  • $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
  • $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
Function INNER-LOOP-E is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices $b$ which are positive-step reachable from the current node $a$ of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from $a$ in case that it visits a vertex $b$ such that $b.keeps_{E}=\mathrm{false}$. It checks whether for each vertex $b$ which are positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{E}=\mathrm{true}$. INNER-LOOP-E decides the following formula:
For each vertex $b\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{E}=\mathrm{true}$.
Function INNER-LOOP-F is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices $b$ which are positive-step reachable from the current node $a$ of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from $a$ in case that in case that it visits a vertex $b$ such that $b.keeps_{F}=\mathrm{false}$. It checks whether for each vertex $b$ which are positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{F}=\mathrm{true}$. INNER-LOOP-F decides the following formula:
For each vertex $b\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node $a$ of the outer loop, $b.keeps_{F}=\mathrm{true}$.
@@
var $a:C\langle E\cup F,\emptyset\rangle$;
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
@@
$a.visited\_outer\_loop:=\mathrm{false}$
$a.visited\_inner\_loop_{E}:=\mathrm{false}$
$a.visited\_inner\_loop_{F}:=\mathrm{false}$
@@
for each vertex $a\in C\langle E\cup F,\emptyset\rangle$ do
if not $a.visited\_outer\_loop$
@@
$a.visited\_outer\_loop:=\mathrm{true}$
${\bf if}$ not ($a.equal_{E}$ or $a.equal_{F}$)
@@
${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
${\bf if}$ not (($a.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b)$
or ($a.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b))$
@@
function $\mathrm{NPDFS-VISIT}(AUX[E;F],a):\;\mathrm{Boolean}$
if not $a.visited\_outer\_loop$
@@
$a.visited\_outer\_loop:=\mathrm{true}$
${\bf if}$ not ($a.equal_{E}$ or $a.equal_{F}$)
@@
${\bf for}$ ${\bf each}$ vertex $b\in Adj(a)$ do
${\bf if}$ not (($a.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b)$
or ($a.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],b)$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],b))$
@@
function $\mathrm{INNER-LOOP-E}(AUX[E;F],a):\;\mathrm{Boolean}$
$a.visited\_inner\_loop_{E}:=\mathrm{true}$;
if $a.keeps_{E}$
${\bf then}$ for each vertex $b\in Adj(a)$ do
@@
$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
@@
$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
@@
$a.visited\_inner\_loop_{E}:=\mathrm{false}$;
@@
function $\mathrm{INNER-LOOP-F}(AUX[E;F],a):\;\mathrm{Boolean}$
$a.visited\_inner\_loop_{F}:=\mathrm{true}$;
if $a.keeps_{F}$
${\bf then}$ for each vertex $b\in Adj(a)$ do
@@
$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
@@
$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
@@
$a.visited\_inner\_loop_{F}:=\mathrm{false}$;
@@
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. NPDFS outputs true on $AUX[E;F]$ if and only if for each $a\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
@@
    $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$.
@@
    $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $a$ to $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
Proof.
By direct inspection of Definition 12.106. ∎
@@
    Algorithm NPDFS on $AUX[E;F]$ in the outer loop visits once each vertex $a\in C\langle E\cup F,\emptyset\rangle$.
@@
    For each vertex $a\in C\langle E\cup F,\emptyset\rangle$ visited in the outer loop,
    • (a)
      $a.equal_{E}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $b$ to $a$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{E}=\mathrm{true}$, or
@@
      $a.equal_{F}=\mathrm{true}$ and for each $b\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from $b$ to $a$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $b.keeps_{F}=\mathrm{true}$.
@@
    Algorithm NPDFS on $AUX[E;F]$ visits at most once each vertex $a\in C\langle E\cup F,\emptyset\rangle$ in the outer loop.
@@
    There exists a vertex $a\in C\langle E\cup F,\emptyset\rangle$ such that Algorithm NPDFS on $AUX[E;F]$ visits $a\in C\langle E\cup F,\emptyset\rangle$ in the outer loop, and Conditions 2.(a) and 2.(b) are false for $a$.
@@
We now continue Examples 3.18, 5.51, 7.72, and 10.96. We run Algorithm NPDFS on $AUX[E;F]$. Then Algorithm NPDFS outputs false. By the last lines of Example 10.96, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Thus Algorithm NPDFS outputs the answer to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
Recall that $n=size(E)+size(F)$. By Statement 4.27, $|C\langle E\cup F,\emptyset\rangle|\leq n$. By Statement 6.63, $|A[E;F]|\leq n$. The running time of depth first search on $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [8, Section 22.3]. Consequently, the running time of depth first search on $AUX[E;F]$ is $O(n)$.
@@
Function INNER-LOOP-E on $AUX[E;F]$ and any vertex $a$ implements a prefix of a depth first search on $AUX[E;F]$ starting from $a$. Accordingly the running time of function INNER-LOOP-E is $O(n)$. Similarly, the running time of function INNER-LOOP-F is $O(n)$. By Statement 4.27, $|C\langle E\cup F,\emptyset\rangle|\leq n$, therefore INNER-LOOP-E and INNER-LOOP-F run at most $n$ times. Accordingly, Algorithm NPDFS runs on $AUX[E;F]$ in $O(n^{2})$ time. ∎

### v2

@@ 12 Main Case 1
$t_{1}\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{d}_{1}$,
$\sigma(t_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma({d}_{1})\mathop{\rightarrow}_{R\langle E,F\rangle}{d}$ for some ${d}\in C\langle E,F\rangle$, where the rewrite rule $\sigma({d}_{1})\dot{\rightarrow}{d}$ is in $R\langle E,F\rangle$, and
$\delta[\sigma(t_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[\sigma({d}_{1})]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\delta[{d}]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
Case 2: $m=0$. Then we consider the rewrite rule $\sigma\dot{\rightarrow}{c}\in R\langle{E\cup F},\emptyset\rangle$. Then we have
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}\delta[{c}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle E\cup F,\emptyset\rangle}$.
@@
$\sigma\mathop{\rightarrow}_{R\langle E,F\rangle}{d}$ for some ${d}\in C\langle E,F\rangle$, where the rewrite rule $\sigma\dot{\rightarrow}{d}$ is in $R\langle E,F\rangle$, and
$\delta[\sigma]\mathop{\rightarrow}_{R\langle E,F\rangle}\delta[{d}]\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
$(\Leftarrow)$ Let $t\in W$ and $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$. Then $s\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. We assumed that GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Consequently, $R\langle E,F\rangle$ or $R\langle F,E\rangle$ simulate $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. Assume that $R\langle E,F\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$. The case $R\langle F,E\rangle$ simulates $R\langle E\cup F,\emptyset\rangle$ on the constant $t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ is similar. By Lemma 12.114, there exists ${e}\in C\langle E,F\rangle$ with ${e}\subseteq t/_{\Theta\langle{E\cup F},\emptyset\rangle}$ such that $s\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}{e}$. By Proposition 4.39, ${e}=t/_{\Theta\langle E,F\rangle}$. For this reason $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. Consequently, for any $s\in t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}$, we have $s\in t/_{\mathop{\leftrightarrow}^{*}_{E}}$. As a consequence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}\subseteq t/_{\mathop{\leftrightarrow}^{*}_{E}}$, and hence $t/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=t/_{\mathop{\leftrightarrow}^{*}_{E}}$. ∎
@@ 12.1 Nested Partial Depth First Search on the Auxiliary dpwpa $AUX[E;F]$ for the GTESs $E$ and $F$
In this way, algorithm NPDFS decides the following formula: For each ${b}\in C\langle E\cup F,\emptyset\rangle$,
  • 1.
    ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$, or
  • 2.
    ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
Function INNER-LOOP-E is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices c{c} which are positive-step reachable from the current node b{b} of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from b{b} in case that it visits a vertex c{c} such that ${c}.keeps_{E}=\mathrm{false}$. It checks whether for each vertex c{c} which are positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{E}=\mathrm{true}$. INNER-LOOP-E decides the following formula:
For each vertex ${c}\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{E}=\mathrm{true}$.
Function INNER-LOOP-F is an inner loop inside the body of the outer loop, it also implements a prefix of a depth first search on those vertices c{c} which are positive-step reachable from the current node b{b} of the outer loop, it can terminate before visiting all vertices of $AUX[E;F]$ which are positive-step reachable from b{b} in case that in case that it visits a vertex c{c} such that ${c}.keeps_{F}=\mathrm{false}$. It checks whether for each vertex c{c} which are positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{F}=\mathrm{true}$. INNER-LOOP-F decides the following formula:
For each vertex ${c}\in C\langle E\cup F,\emptyset\rangle$ which is positive-step reachable from the current node b{b} of the outer loop, ${c}.keeps_{F}=\mathrm{true}$.
@@
var ${b}:C\langle E\cup F,\emptyset\rangle$;
for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
@@
${b}.visited\_outer\_loop:=\mathrm{false}$
${b}.visited\_inner\_loop_{E}:=\mathrm{false}$
${b}.visited\_inner\_loop_{F}:=\mathrm{false}$
@@
for each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ do
if not b.visited_outer_loop{b}.visited\_outer\_loop
@@
${b}.visited\_outer\_loop:=\mathrm{true}$
$\bf if$} not ( ${b}.equal_{E}$ or ${b}.equal_{F}$)
@@
$\bf for$} $\bf each$} vertex ${c}\in Adj({b})$ do
$\bf if$} not $($${b}.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c})$
or (${b}.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c}))$
@@
function $\mathrm{NPDFS-VISIT}(AUX[E;F],{b}):\;\mathrm{Boolean}$
if not b.visited_outer_loop{b}.visited\_outer\_loop
@@
${b}.visited\_outer\_loop:=\mathrm{true}$
$\bf if$} not (${b}.equal_{E}$ or ${b}.equal_{F}$)
@@
$\bf for$} $\bf each$} vertex ${c}\in Adj({b})$ do
$\bf if$} not $($${b}.equal_{E}$ and $\mathrm{INNER-LOOP-E}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c})$
or (${b}.equal_{F}$ and $\mathrm{INNER-LOOP-F}(AUX[E;F],{c})$ and $(\mathrm{NPDFS-VISIT}(AUX[E;F],{c}))$
@@
function $\mathrm{INNER-LOOP-E}(AUX[E;F],{b}):\;\mathrm{Boolean}$
${b}.visited\_inner\_loop_{E}:=\mathrm{true}$;
if ${b}.keeps_{E}$
$\bf then$} for each vertex ${c}\in Adj({b})$ do
@@
${b}.visited\_inner\_loop_{E}:=\mathrm{false}$;
@@
${b}.visited\_inner\_loop_{E}:=\mathrm{false}$;
@@
${b}.visited\_inner\_loop_{E}:=\mathrm{false}$;
@@
function $\mathrm{INNER-LOOP-F}(AUX[E;F],{b}):\;\mathrm{Boolean}$
${b}.visited\_inner\_loop_{F}:=\mathrm{true}$;
if ${b}.keeps_{F}$
$\bf then$} for each vertex ${c}\in Adj({b})$ do
@@
${b}.visited\_inner\_loop_{F}:=\mathrm{false}$;
@@
${b}.visited\_inner\_loop_{F}:=\mathrm{false}$;
@@
${b}.visited\_inner\_loop_{F}:=\mathrm{false}$;
@@
Let $E$ and $F$ be GTESs over a unary signature $\Sigma$. NPDFS outputs true on $AUX[E;F]$ if and only if for each ${b}\in C\langle E\cup F,\emptyset\rangle$, Conditions 1 or 2 hold:
@@
    ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$.
@@
    ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from b{b} to c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
We get the following lemma by direct inspection of Definition 12.116.
@@
    Algorithm NPDFS on $AUX[E;F]$ in the outer loop visits once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$.
@@
    For each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ visited in the outer loop,
@@
      ${b}.equal_{E}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from c{c} to b{b} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{E}=\mathrm{true}$, or
    • (c)
      ${b}.equal_{F}=\mathrm{true}$ and for each ${c}\in C\langle E\cup F,\emptyset\rangle$ if there is a path of positive length or a cycle from c{c} to b{b} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then ${c}.keeps_{F}=\mathrm{true}$.
@@
    Algorithm NPDFS on $AUX[E;F]$ visits at most once each vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ in the outer loop.
@@
    There exists a vertex ${b}\in C\langle E\cup F,\emptyset\rangle$ such that Algorithm NPDFS on $AUX[E;F]$ visits ${b}\in C\langle E\cup F,\emptyset\rangle$ in the outer loop, and Conditions 2.(a) and 2.(b) are false for b{b}.
@@
We now continue Examples 3.19, 5.52, 7.73, and 10.106. We run Algorithm NPDFS on $AUX[E;F]$. Then Algorithm NPDFS outputs false. By the last lines of Example 10.106, GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together do not simulate the GTRS $R\langle E\cup F,\emptyset\rangle$. Thus Algorithm NPDFS outputs the answer to the question whether GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ together simulate the GTRS $R\langle E\cup F,\emptyset\rangle$.
@@
Recall that $n=\mathrm{size}(E)+\mathrm{size}(F)$. By Statement 4.28, $|C\langle E\cup F,\emptyset\rangle|\leq n$. By Statement 6.64, $|A[E;F]|\leq 2n$. The running time of depth first search on $AUX[E;F]$ is $O(|C\langle E\cup F,\emptyset\rangle|+|A[E;F]|)$ [13, Section 22.3]. Consequently, the running time of depth first search on $AUX[E;F]$ is $O(n)$.
@@
Function INNER-LOOP-E on $AUX[E;F]$ and any vertex b{b} implements a prefix of a depth first search on $AUX[E;F]$ starting from b{b}. Accordingly the running time of function INNER-LOOP-E is $O(n)$. Similarly, the running time of function INNER-LOOP-F is $O(n)$. By Statement 4.28, $|C\langle E\cup F,\emptyset\rangle|\leq n$, therefore INNER-LOOP-E and INNER-LOOP-F run at most $n$ times. Accordingly, Algorithm NPDFS runs on $AUX[E;F]$ in $O(n^{2})$ time. ∎